OpenStudy (anonymous):

The temperature change, T, in a patient generated by a dose, D, of a drug is given by T=(C/2-D/3)D^2 where C is a postitive constant what dosage maximizes the temperature change? the sensitivity of the body, at dosage D, to the drug is defined as dT/dD what dosage maximizes sensitivity.

OpenStudy (anonymous):

the function is \[T(D)=\frac{\frac{C}{2}-\frac{D}{3}}{D^2}\]

OpenStudy (anonymous):

C is a constant in this yes? do take derivative wrt D , set = 0 and solve for D

OpenStudy (anonymous):

\[T=(\frac{C}{2}-\frac{D}{3})D^2\]

OpenStudy (anonymous):

oh wait not what you wrote. is it \[T(D)=(\frac{C}{2}-\frac{D}{3})D^2\]?

OpenStudy (anonymous):

got it. multiply out to get \[\frac{CD^2}{2}-\frac{D^3}{3}\]

OpenStudy (anonymous):

take the derivative wrt D to get \[T'(D)=CD-D^2\]

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

set = 0 and solve for D via \[CD-D^2=0\] \[D(C-D)=0\] \[D=0\] (unlikely) or \[D=C\] which is your answer.

OpenStudy (anonymous):

i have both

OpenStudy (anonymous):

well it is certainly unlikely that D = 0 is an answer since if D = 0 you get T(D) = 0

OpenStudy (anonymous):

there are no values in the question so i would say D=C

OpenStudy (anonymous):

so what's that? for the dosage that maximizes the temperature change if so how do you find part II